LaTEX Reference Sheet

for LaTEX 2.09 on Athena

ó \'{o}  ö \"{o}  ǒ \v{o}  ô \^{o}  oo \t{oo}  ō \={o}  ȯ \.{o}
ò \`{o}  ő \H{o}  õ \~{o}  ŏ \u{o}  o̧ \c{o}      ō \b{o}      o. \d{o}
       œ \oe  æ \ae  å \aa  ø \o  ł \l  ¿ ?`
      Œ \OE  Æ \AE  Å \AA  Ø \O  Ł \L  ¡ !`
       † \dag  ‡ \ddag  § \S  ¶ \P  £ \pounds    \begin{matrix}c \\ ◯ \end{matrix} \copyright
              x^{y−1} x^{y-1}       ^{6} √ \sqrt[6]{. . . } ∑ \sum  ∫ \int
               x′ x'            \frac{a}{b} \frac\{a\}\{b\} ∏ \prod  ∮ \oint

\begin{matrix}c \\ ◯ \end{matrix} 1996 Student Information Processing Board of the Massachusetts Institute of Technology. For more information about L^{A}TEX, pick up a copy of Inessential LaTEX in the SIPB office, room w20-557.

          This document is $Revision: 0.0 $. Last modified $Date: 1996/10/07 21:53:32 $.